OpenStudy (anonymous):

Paras custom tee experienced fixed cost of $300 and a variable cost $4 a shift write and equation that can be used to determine the total expense encountered by paras custom tee. The cost of producing 7 shirts is

pooja195 (pooja195):

C(x) = 300 + 4x x be the number of shirts and let C(x) be the total cost of producing x shirts. You want to find the cost of producing 7 shirts right? so plug in 7 in place of x like this \[\huge~\rm~C(x)= 300 + 4(7)\] solve

OpenStudy (anonymous):

328

pooja195 (pooja195):

Good :) so $328.00

pooja195 (pooja195):

And then i gave you the equation above.

OpenStudy (anonymous):

what is "x" if (x-3)^2=x^2+(-3)^2

imqwerty (imqwerty):

i will help u and i won't take owlbucks :) \[(x-3)^2=x^2+(-3)^2 => (x-3)^2-(x)^2 =(-3)^2\] apply this identity\[a^2-b^2=(a+b)(a-b)\] \[(x-3-x)(x-3+x)=9\] \[(-3)(2x-3)=9\]divide both sides by -3 \[(2x-3)=-3 =>2x=6 =>x=3\] :)